CH/DOC/94/09

Authorship

Practice for papers from ANU

Author: B.R. Fulton

Date: 21 April 1994

Practice for authors on papers

Last year we discussed the question of authorship on papers from
the first period of running at ANU and agreed that those involved
in the installation and commissioning periods would also be included
on the author list in addition to those involved in the actual runs.
This was an obvious acknowledgement of the work done by the first
groups out there, withoutwhose hard work none of the experiments
could have taken place.

I have been asked whether this policy still applies and hence have
placed this item on the agenda.  It is my own feeling that for the
running period just finished, and subsequent periods, we should
revert to a more normal approach.  This would mean that the author
list would reflect those who participated in the run plus any others
who contributed to a significant degree during the analysis or
interpretation stages.  This would reflect our previous practice at
the NSF and also be in line with other groups.  It is of course true
that in the pattern of operation where we pack a number of runs
into one period at ANU, the success of a particular run might depend
on earlier work done by an individual who has returned before the
run starts.  I would hope that such cases would be dealt with as
an when they arise.
Perhaps this could be discussed at the meeting and if agreement is
reached we can record the policy.
Brian
21.4.94